# Effective Whitney Stratification of Real Algebraic Varieties

### v3

@@ 6.2. Comparisons, Practical Considerations and Tests
Below we give a table comparing the polar variety based Whitney stratification methods to the other methods which are capable of computing Whitney stratifications in practice (all of which are based on the conormal stratification procedure reviewed in Section 2.1). We note that the method of [32] is not included, as it was unable to stratify even the Whitney umbrella, which is the simplest possible example, within 24 hours. Similarly, using the QuantifierElimination package in Maple we implemented the algorithm of [42, 45] for the Whitney umbrella. The code ran for slightly longer than 24 hours, using over 50 GB of RAM (on a workstation with an Intel Xeon W-3365 processor and 1000 GB of RAM), before crashing with a "Stack Limit Reached" error, even if we allowed unlimited stack size. In contrast, for the Whitney umbrella example, both the conormal-based methods reviewed in Section 2.1 and the method based on the polar variety computation from Section 3 executed successfully in less than 0.1 seconds (in Macaulay2 and working over the rationals, $\mathbb{Q}$) using a barely measurable amount of RAM.
Input Assoc. Primes [26] Block Order [25] WhitneyPolar
Whitney Umbrella 0.06s 0.05s 0.05s
Whitney Cusp 0.11s 0.07s 0.06s
$X_{1}$, see (12) 2.1s 1.8s 0.2s
$X_{2}$, see (13) 20.0s 2.0s
$X_{3}$, see (14) 1.4s 0.8s 0.2s
$X_{4}$, see (15) 5.2s 5.2s 0.1s
Smooth 0.05s 0.05s 0.05s
Barth Sextic 0.9s 0.9s 0.9s
Table 1. Run times of several examples; all computations were run in Macaulay2 version 1.24.11 on a laptop with an Intel i7-1185G7 processor and 31 GB of RAM. For more on the computational setup see Remark 6.6. Computations stopped after 8 hours are denoted –.
Remark 6.6.
All computations in Table 1 are run over a finite field of prime characteristic, 32749, rather than, e.g. ,over the rationals. The reason for this is that the practical performance of the WhitneyPolar algorithm is dependent on the availability of an implementation of a Gröbner basis algorithm such as F4 [16] which works well for dense systems in practice. Currently, in Macaulay2, the F4 algorithm can only be reliability used over a finite field. Additionally, implementations in, e.g., Julia, where a reliable F4 Gröbner basis algorithm over $\mathbb{Q}$ is already available, would be expected to match the relative treads in the table. All algorithms above also execute correctly over $\mathbb{Q}$; however, the need of the WhitneyPolar algorithm to fall back on a Buchberger type Gröbner basis in Macaulay2 makes the comparison less informative.
Below we give the equations for the examples presented in Table 1. Examples $X_{1}$ through $X_{4}$ are taken from [25] and were chosen as they have a non-trivial Whitney stratification in the sense that the strata are not obtained by simply computing iterated singular loci. These exhibit varying levels of complexity in terms of degree of generators, number of variables, density of polynomials, etc. The other varieties, besides the Whitney cusp and umbrella, have Whitney stratifications given by iterated singular loci; these have been included as control examples. Of these, the smooth example is obtained by intersecting 2 random quadrics in 10 variables. While the Barth Sextic is naturally written in 3 variables, we require an additional variable to represent the golden ratio $\frac{\sqrt{5}+1}{2}$ algebraically. Thus, our version of the sextic is represented by an ideal of degree 12 in 4 variables.
@@
$\begin{split}X_{2}=\mathbf{V}(&x_{1}^{2}x_{2}^{2}x_{3}-2x_{1}^{2}x_{2}x_{3}x_{4}-2x_{1}^{2}x_{2}x_{3}x_{5}+x_{1}^{2}x_{3}x_{4}^{2}+2x_{1}^{2}x_{3}x_{4}x_{5}+x_{1}^{2}x_{3}x_{5}^{2}-2x_{1}x_{2}^{3}x_{3}+\\ &4x_{1}x_{2}^{2}x_{3}x_{4}+4x_{1}x_{2}^{2}x_{3}x_{5}-x_{1}x_{2}^{2}x_{4}-2x_{1}x_{2}x_{3}x_{4}^{2}-6x_{1}x_{2}x_{3}x_{4}x_{5}-2x_{1}x_{2}x_{3}x_{5}^{2}+\\ &x_{1}x_{2}x_{4}^{2}+x_{1}x_{2}x_{4}x_{5}+2x_{1}x_{3}x_{4}^{2}x_{5}+2x_{1}x_{3}x_{4}x_{5}^{2}-x_{1}x_{4}^{2}x_{5}+x_{2}^{4}x_{3}-2x_{2}^{3}x_{3}x_{4}-\\ &2x_{2}^{3}x_{3}x_{5}+x_{2}^{3}x_{4}+x_{2}^{2}x_{3}x_{4}^{2}+4x_{2}^{2}x_{3}x_{4}x_{5}+x_{2}^{2}x_{3}x_{5}^{2}-2x_{2}^{2}x_{4}^{2}-x_{2}^{2}x_{4}x_{5}-2x_{2}x_{3}x_{4}^{2}x_{5}-\\ &2x_{2}x_{3}x_{4}x_{5}^{2}+x_{2}x_{4}^{3}+2x_{2}x_{4}^{2}x_{5}+x_{3}x_{4}^{2}x_{5}^{2}-x_{4}^{3}x_{5})\subset\mathbb{C}^{5}\end{split}$ (13)
$X_{3}=\mathbf{V}(x_{1}^{3}-2x_{1}^{2}x_{4}+x_{1}^{3}+x_{5}^{2}x_{4}+x_{2}^{2}x_{1}-x_{2}x_{1}x_{3}+x_{5}^{3})\subset\mathbb{C}^{5}$ (14)
$X_{4}=\mathbf{V}\left(x_{1}^{2}x_{3}-x_{2}^{2},\,x_{2}^{4}-x_{1}x_{2}^{2}-x_{3}x_{4}^{2},\,x_{1}^{2}x_{2}^{2}-x_{1}^{3}-x_{4}^{2}\right)\subset\mathbb{C}^{4}$ (15)

### v4

@@ 6.1. Other Methods
We have compared WhitneyPolar to all known methods for computing practical Whitney stratifications in practice; all of these are based on the conormal stratification procedure reviewed in Section˜2.1. This comparison occupies the first three columns of Table 1. For reference, we also give the run time for computing a Cylindrical Algebraic Decomposition (CAD) for each given variety. Conversely, we do not include the method of [30] in our results because it was unable to stratify the Whitney umbrella – which is the simplest nontrivial input – after running for 24 hours. We also implemented the algorithm of [38, 41] for the Whitney umbrella, using the standard QuantifierElimination package in Maple. This code ran for slightly longer than 24 hours, using over 50 GB of RAM (on a workstation with an Intel Xeon W-3365 processor and 1000 GB of RAM). It ultimately crashed with a "Stack Limit Reached" error, even if we allowed unlimited stack size.
6.2. Sample Code
An example of the snippet of code which is run in Macaulay2 [18] applied to example $X_{1}$ from Table 1 is given below [^9] Note that running this code successfully requires Macaulay2 version 1.25.06 or higher and version 2.23 or higher of the WhitneyStratifications [22] package.:
needsPackage"WhitneyStratifications"
R=QQ[x_1..x_4]
I=ideal(x_1^6+x_2^6+x_1^4*x_3*x_4+x_3^3)
elapsedTime V=whitneyStratify(I)
elapsedTime W=whitneyStratify(I,AssocPrimes=>false)
elapsedTime Vpol=whitneyStratifyPol(I,Algorithm=>"msolve")
In the code snippet above the stratification V is computed using the method of [24], the stratification W is computed using the method of [23], and the stratification Vpol is computed using the WhitneyPolar algorithm introduced above in Section 3.
6.3. Gröbner Basis Computation
For all Whitney stratification algorithms demonstrated in Table 1, we have provided the run time both over $\mathbb{F}_{32749}$, a finite field of characteristic 32749, and over the rationals. In practice, efficient implementations of Gröbner basis algorithms over $\mathbb{Q}$ often first compute over different large finite fields, and then employ rational reconstruction (this is used in [3] for instance). A similar multi-modular philosophy could be applied to Whitney stratification computation as a whole. On larger examples, the timings in the table seem to indicate there could be a significant benefit to doing so.
6.4. Hardware
All computations in the first three columns of Table 1 were run using Version 2.23 of the WhitneyStratifications [22] package in Macaulay2 [18] version 1.25.06. The last column used the CylindricalDecomposition command in Mathematica and was run with Wolfram 14.2.1. All computations were carried out on a laptop with an Intel Ultra 7 processor (258V) and 32 GB of RAM. The WhitneyPolar algorithm uses the msolve [3] Gröbner basis library (version 0.9.0) via the Msolve Macaulay2 package. Computations which used greater than 32GB we also run again on a desktop with 64GB of RAM available; in all cases these computations again exhausted all available RAM before the end of the 11 hour run. On all runs, WhitneyPolar used no more that 400MB of memory.
6.5. The Input Varieties
Below we give the equations for the examples presented in Table 1. The Whitney Cusp and Umbrella are the simplest examples of singular varieties whose Whitney stratification can not be obtained by computing iterated singular loci. Examples $X_{1}$ through $X_{4}$ are taken from [23]; these are more complicated and also have non-trivial Whitney stratifications.
@@ 6.5. The Input Varieties
$X_{2}=\mathbf{V}\left(x_{1}^{2}x_{3}-x_{2}^{2},\,x_{2}^{4}-x_{1}x_{2}^{2}-x_{3}x_{4}^{2},\,x_{1}^{2}x_{2}^{2}-x_{1}^{3}-x_{4}^{2}\right)\subset\mathbb{C}^{4}$ (14)
$X_{3}=\mathbf{V}\left((x_{1}^{3}-2x_{1}^{2}x_{4}+x_{1}^{3}+x_{5}^{2}x_{4}+x_{2}^{2}x_{1}-x_{2}x_{1}x_{3}+x_{5}^{3}\right)\subset\mathbb{C}^{5}$ (15)
$\begin{split}X_{4}=\mathbf{V}\Big{(}&x_{1}^{2}x_{2}^{2}x_{3}-2x_{1}^{2}x_{2}x_{3}x_{4}-2x_{1}^{2}x_{2}x_{3}x_{5}+x_{1}^{2}x_{3}x_{4}^{2}+2x_{1}^{2}x_{3}x_{4}x_{5}+x_{1}^{2}x_{3}x_{5}^{2}-2x_{1}x_{2}^{3}x_{3}+\\ &4x_{1}x_{2}^{2}x_{3}x_{4}+4x_{1}x_{2}^{2}x_{3}x_{5}-x_{1}x_{2}^{2}x_{4}-2x_{1}x_{2}x_{3}x_{4}^{2}-6x_{1}x_{2}x_{3}x_{4}x_{5}-2x_{1}x_{2}x_{3}x_{5}^{2}+\\ &x_{1}x_{2}x_{4}^{2}+x_{1}x_{2}x_{4}x_{5}+2x_{1}x_{3}x_{4}^{2}x_{5}+2x_{1}x_{3}x_{4}x_{5}^{2}-x_{1}x_{4}^{2}x_{5}+x_{2}^{4}x_{3}-2x_{2}^{3}x_{3}x_{4}-\\ &2x_{2}^{3}x_{3}x_{5}+x_{2}^{3}x_{4}+x_{2}^{2}x_{3}x_{4}^{2}+4x_{2}^{2}x_{3}x_{4}x_{5}+x_{2}^{2}x_{3}x_{5}^{2}-2x_{2}^{2}x_{4}^{2}-x_{2}^{2}x_{4}x_{5}-2x_{2}x_{3}x_{4}^{2}x_{5}-\\ &2x_{2}x_{3}x_{4}x_{5}^{2}+x_{2}x_{4}^{3}+2x_{2}x_{4}^{2}x_{5}+x_{3}x_{4}^{2}x_{5}^{2}-x_{4}^{3}x_{5}\Big{)}\subset\mathbb{C}^{5}\end{split}$ (16)
Examples $X_{5}$ and $X_{6}$ are also constructed to have non-trivial Whitney stratifications.
$X_{5}=\mathbf{V}\left(x_{6}x_{7}^{2}+x_{7}^{3}+x_{6}^{2}-x_{1}x_{7}-x_{7}^{2}+x_{2},\,x_{1}^{2}x_{2}^{2}-x_{1}^{3}+x_{1}x_{7}^{2}-x_{3}x_{7}^{2}-x_{5}^{2}\right)\subset\mathbb{C}^{7}$ (17)
$\begin{split}X_{6}=\mathbf{V}\Big{(}&3\,x_{2}x_{3}x_{5}+2\,x_{1}x_{4}x_{5}-x_{2}x_{4}x_{5}-x_{3}x_{4}x_{5}-2\,x_{1}x_{5}x_{6}-5\,x_{3}x_{5}x_{6}+2\,x_{4}x_{5}x_{6},\\ &x_{2}x_{4}^{2}-x_{1}x_{4}x_{5}-3\,x_{2}x_{3}x_{6}+x_{1}x_{5}x_{6}+3\,x_{3}x_{5}x_{6}-x_{4}x_{5}x_{6},\\ &\,x_{2}^{2}x_{4}-x_{1}x_{2}x_{5}+2\,x_{1}x_{2}x_{6}-x_{2}x_{3}x_{6}-2\,x_{2}x_{4}x_{6}+x_{3}x_{5}x_{6},\\ &\,3\,x_{2}^{2}x_{3}+2\,x_{1}x_{2}x_{4}-x_{2}x_{3}x_{4}-x_{1}x_{2}x_{5}-6\,x_{2}x_{3}x_{6}+x_{3}x_{5}x_{6}\Big{)}\subset\mathbb{C}^{6}\end{split}$ (18)
These examples exhibit varying levels of complexity in terms of degree of generators, number of variables, density of polynomials, etc. Our next example $X_{7}$ is the determinantal variety defined by the $3\times 3$ minors of the matrix
[…]
Explicitly, we have
$\begin{split}X_{7}=\mathbf{V}(&x_{1}x_{4}x_{5}-x_{3}x_{5}x_{6}+x_{4}x_{5}x_{6}-x_{4}^{2}x_{7}-x_{1}x_{5}x_{7}+x_{3}x_{7}^{2},\\ &x_{1}^{2}x_{5}+x_{2}x_{4}x_{6}-x_{3}x_{5}x_{6}-x_{1}x_{2}x_{7}-x_{1}x_{4}x_{7}+x_{3}x_{7}^{2},\\ &x_{2}x_{4}^{2}+x_{1}x_{3}x_{5}-x_{3}x_{4}x_{5}-x_{3}x_{5}x_{6}+x_{4}x_{5}x_{6}-x_{2}x_{3}x_{7}-x_{4}^{2}x_{7}-x_{1}x_{5}x_{7}+x_{3}x_{5}x_{7}+x_{3}x_{7}^{2},\\ &x_{1}x_{2}x_{4}-x_{2}x_{3}x_{6}+x_{2}x_{4}x_{6}-x_{1}x_{2}x_{7}+x_{1}x_{3}x_{7}-x_{1}x_{4}x_{7}-x_{3}x_{4}x_{7}+x_{3}x_{7}^{2})\subset\mathbb{C}^{7}\end{split}$ (19)
The last two varieties in Table 1 have Whitney stratifications given by iterated singular loci; these have been included as control examples. The first of these is smooth and has been obtained by intersecting 2 random quadrics in 10 variables. The second one, called the Barth Sextic is naturally written in 3 variables, however we require an additional variable to represent the golden ratio $\frac{\sqrt{5}+1}{2}$ algebraically. Thus, our version of the sextic is represented by an ideal of degree 12 in 4 variables.
Input Assoc. Primes [24] Block Order [23] WhitneyPolar CAD
$\mathbb{F}_{32749}$ $\mathbb{Q}$ $\mathbb{F}_{32749}$ $\mathbb{Q}$ $\mathbb{F}_{32749}$ $\mathbb{Q}$
Whitney Umbrella 0.07s 0.07s 0.05s 0.05s 0.06s 0.06s 0.4s
Whitney Cusp 0.09s 0.09s 0.07s 0.07s 0.06s 0.06s 0.4s
$X_{1}$, see (13) 4.8s 2.9s 4.5s 2.7s 0.15s 2.5s 0.4s
$X_{2}$, see (14) 9.5s 58.9s 9.2s 58.3s 0.1s 0.2s 0.4s
$X_{3}$, see (15) 1.1s 1.6s 0.6s 0.9s 0.2s 0.3s 0.9s
$X_{4}$, see (16) $m$ $m$ 20.7s 747.2s 1.9s 3.4s $m$
$X_{5}$, see (17) $m$ $m$ $m$ $m$ 1.6s 77.2s $t$
$X_{6}$, see (18) $t$ $t$ $t$ $t$ 1.4s 3.9s $m$
$X_{7}$, see (19) $t$ $t$ $t$ $t$ 52.5s 208.8s $m$
Smooth 0.08s 0.4s 0.08s 0.4s 0.08s 0.4s $t$
Barth Sextic 0.6s 0.6s 0.6s 0.6s 0.6s 0.6s 12717.7s
Table 1. The first three columns show run times of various practical Whitney stratification algorithms on several example varieties; the run time for Cylindrical Algebraic Decomposition (CAD) computation is also included in the last column. Computations which failed to terminate after 11 hours are denoted $t$. Computations which exhausted 64GB of memory and terminated are denoted $m$. For each run we highlight the fastest algorithm in bold.
6.6. The Results
Regarding the experiments in Table 1, we note that the practical benefits of the new WhitneyPolar algorithm become particularly apparent on larger examples with denser polynomials, e.g. $X_{4}$, $X_{5}$, $X_{6}$ and $X_{7}$. This is true both in comparison to the other Whitney stratification methods from [23, 24], as well as when compared to alternative methods to study systems of real polynomials such as Mathematica’s highly refined CAD implementation.